# The Unbounded Denominators Conjecture

### v3

@@ 2.4.1. Lemma on the lexicographically lowest coefficient
Consider $G(\mathbf{z})\in\mathbf{C}\llbracket\mathbf{z}\rrbracket\smallsetminus\{0\}$ holomorphic on the closed unit polydisc $\{\mathbf{z}\mid\max_{i=1}^{d}|z_{i}|\leq 1\}$, and let $c\,\mathbf{z^{n}}$ be the lexicographically minimal monomial. Then
@@
By assumption, the lexicographically minimal monomial in $H(0,\mathbf{z}^{\prime})\in\mathbf{C}\llbracket\mathbf{z}^{\prime}\rrbracket$ is $c\,\mathbf{z}^{\prime\mathbf{n}^{\prime}}$, where $\mathbf{n}=(n_{1},\mathbf{n}^{\prime})$. Therefore the induction hypothesis gives
@@
Thus, with $c\,\mathbf{z^{n}}$ the lexicographically lowest monomial in $G(\mathbf{z})$, we get from (2.4.2) and Lemma 2.1.2 our Cauchy upper bound:
$\log{|c|}\leq\int_{\mathbf{T}^{d}}\log{|F(\varphi(z_{1}),\ldots,\varphi(z_{d}))|}\,\mu_{\mathrm{Haar}}$
(2.4.6) $\leq dD\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}+\kappa C\,\alpha+o(\alpha),$
@@
on the split-variables constituents in (2.1.3). By our normalizations $x(t)\in t+t^{2}\mathbf{Q}\llbracket t\rrbracket$ and $\varphi(z)\in\varphi^{\prime}(0)z+z^{2}\mathbf{C}\llbracket z\rrbracket$, the lexicographically lowest term of $G(\mathbf{z})\in\mathbf{C}\llbracket\mathbf{z}\rrbracket$ is equal to $\varphi^{\prime}(0)^{\beta}$ times the lexicographically lowest term of $F(x(\mathbf{t}))\in\mathbf{Z}\llbracket\mathbf{t}\rrbracket$, where $\beta:=|\mathbf{n}|=n_{1}+\cdots+n_{d}\geq\alpha$ is the common total degree of these lexicographically lowest terms in $G(\mathbf{z})$ and $F(x(\mathbf{t}))$. By (2.4.7), this entails that the nonzero coefficient
@@ 2.5. A second alternative proof
The remainder of §2 presents our original argument for Theorem 2.0.1, with the thought that it could still be useful for other settings including potential theory (see 2.5.4). Like § 2.2 and unlike § 2.4, it is based on the leading order jet rather than the overall lexicographically lowest monomial in $F(\mathbf{x})$, and on the pointwise Cauchy integral formula instead of on plurisubharmonicity. Contrastingly to both, it employs a cross-variables equidistribution idea.
@@ 2.5.1. Equidistribution
We start out the same way with Lemma 2.1.2, but now aim to extrapolate based directly on the pointwise Cauchy bound. The key idea here is that upon substituting $x_{j}=\varphi(z_{j})$ into (2.1.3), the $d\to\infty$ equidistribution on the circle of the uniform independent and identically distributed points $z_{1},\ldots,z_{d}$ will normally get the constituent monomials in (2.1.3) to grow at most at the integrated exponential rate of $dD\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}$. The problem with directly applying the Cauchy bound as in [And89, VIII 1.6] is that it involves a pointwise upper bound on the intervening functions $|p(\varphi(\mathbf{z}))^{\mathbf{k}}|$ on the unit polycircle $\mathbf{z}\in\mathbf{T}^{d}$, and while the Monte Carlo heuristic applies on the majority of $\mathbf{T}^{d}$ under $d\to\infty$, with a probability tending to $1$ roughly speaking at a rate exponential in d-d (this follows by Hoeffding’s concentration inequality with (2.5.7) below), the peaks at the biased part of $\mathbf{T}^{d}$ get overwhelmingly large, and a direct extrapolation with (2.1.3) in this way still only leads to a dimension bound with $\sup_{|z|=1}\log{|p\circ\varphi|}$.
@@
The idea for sifting out the equidistributed tuples $(z_{1},\ldots,z_{d})$ is the following. If the points $z_{1},\ldots,z_{d}$ are poorly distributed in the uniform measure of the circle, the quantity $|V(\mathbf{z})|$ is uniformly exponentially small in $d^{2}$ (Lemma 2.5.8 below). This plays off against the $d^{d/2}=\exp(o(d^{2}))$ bound of Lemma 2.5.2 to sift out the equidistributed points in our pointwise upper bound in the Cauchy integral formula when we extrapolate in § 2.4.2 below. Liouville’s Diophantine lower bound still succeeds like in André [And04, §5], thanks to the chain rule and the integrality property of coefficients of (2.5.1), but at the Cauchy upper bound we are now aided by the fact that $V(\mathbf{z})^{M}$ is extremely small (an exponential in $Md^{2}$, see (2.5.9)) at the peaks of the pointwise Cauchy bound, where the point $(z_{1},\ldots,z_{d})$ is poorly distributed, while still not too large (subexponential in $Md^{2}$, thanks to Lemma 2.5.2) uniformly throughout the whole polycircle $\mathbf{T}^{d}$.
@@
(2.5.7) $D(z_{1},\ldots,z_{d})\leq 3\Big(\frac{1}{K+1}+\sum_{k=1}^{K}\frac{1}{k}\Big|\frac{z_{1}^{k}+\cdots+z_{d}^{k}}{d}\Big|\Big)\quad\forall K\in\mathbf{N},$
@@
Thus we introduce another parameter $\varepsilon>0$, which in the end will be let to approach 0 but only after $d\to\infty$, and we divide the points $\mathbf{z}\in\mathbf{T}^{d}$ into two groups according to whether $D(z_{1},\ldots,z_{d})<\varepsilon$ (the well-distributed points) or $D(z_{1}\ldots,z_{d})\geq\varepsilon$ (the poorly distributed points). For the well-distributed group we use the Koksma inequality (2.5.6), and for the poorly distributed group we take advantage of the overwhelming damping force of the Vandermonde factor.
@@
Since the qualitative result suffices for our purposes here, we give a soft proof based on compactness. The following argument borrows from Bombieri and Gubler’s exposition [BG06, page 103] of Bilu’s equidistribution theorem. Suppose to the contrary that there is an $\varepsilon\in(0,1]$ and an infinite sequence $(z_{1}^{(d)},\ldots,z_{d}^{(d)})\in\mathbf{T}^{d}$ such that
(2.5.10) $\lim_{d\to\infty}\frac{1}{\binom{d}{2}}\sum_{1\leq i<j\leq d}\log{\frac{1}{|z_{i}^{(d)}-z_{j}^{(d)}|}}\leq 0,$
@@
(2.5.11) $\forall d,\quad\quad D(z_{1}^{(d)},\ldots,z_{d}^{(d)})\geq\varepsilon.$
@@
across all probability measures $\nu$ supported by $K$. By symmetry, we have $\mu_{\mathbf{T}}=\mu_{\mathrm{Haar}}$, and since $I(\mu_{\mathrm{Haar}})=0$, but $\mu\neq\mu_{\mathrm{Haar}}$, we have the strict inequality
@@ 2.5.2. Damping the Cauchy estimate
We combine Lemmas 2.5.2 and 2.5.8 for our choice of the damping term $V(\mathbf{z})^{M}$. In the following, all asymptotics are taken under $\alpha\to\infty$ with respect to all other parameters.
@@
$D(z_{1},\ldots,z_{d})<\varepsilon\quad\Longrightarrow$
$\log{|F(\varphi(z_{1}),\ldots,\varphi(z_{d}))|}\leq dD\,\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}+\kappa C\alpha+O_{p,\varphi}(\varepsilon\,dD)+o(\alpha).$
@@
On the poorly distributed but exceptional part $D(z_{1},\ldots,z_{d})\geq\varepsilon$, the sum in (2.5.13) can get as large as $d\sup_{|z|=1}\log{|p\circ\varphi|}$. This trivial bound gives
(2.5.14) $\forall\mathbf{z}\in\mathbf{T}^{d},\quad\log{|F(\varphi(z_{1}),\ldots,\varphi(z_{d}))|}\leq dD\sup_{|z|=1}\log^{+}{|p\circ\varphi|}+\kappa C\alpha+o(\alpha).$
@@
for the remainder of the proof of Theorem 2.0.5 (at the end we will firstly take $d\to\infty$, and only then $\varepsilon\to 0$), and we select the Vandermonde exponent
@@
(2.5.18) $\sup_{\mathbf{z}\in\mathbf{T}^{d}:\,D(z_{1},\ldots,z_{d})\leq\varepsilon}\log{|V(\mathbf{z})^{M}F(\varphi(\mathbf{z}))|}$
$\leq dD\,\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}+\kappa C\alpha+O_{p,\varphi}(\varepsilon\alpha)+O_{\varepsilon,p,\varphi}\Big(\frac{\log{d}}{d}\alpha\Big)+o(\alpha).$
@@
(2.5.19) $H(\mathbf{z}):=V(\mathbf{z})^{M}F(\varphi(z_{1}),\ldots,\varphi(z_{d}))=:\sum_{\mathbf{n}\in\mathbf{N}_{0}^{d}}c(\mathbf{n})\,\mathbf{z^{n}}\in\mathbf{C}\llbracket\mathbf{z}\rrbracket,$
convergent on the closed unit disc $\|\mathbf{z}\|\leq 1$. For each $\mathbf{n}\in\mathbf{N}_{0}^{d}$, the $\mathbf{z^{n}}$ coefficient of $H(\mathbf{z})$ is given by the Cauchy integral formula
@@
(2.5.21) $\forall\mathbf{n}\in\mathbf{N}_{0}^{d},\quad\quad|c(\mathbf{n})|\leq\sup_{\mathbf{z}\in\mathbf{T}^{d}}|H(\mathbf{z})|.$
@@
(2.5.22) $\log{|c(\mathbf{n})|}\leq dD\,\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}$
$+O(\kappa\alpha)+O_{p,\varphi}(\varepsilon\,\alpha)+O_{\varepsilon,p,\varphi}\Big(\frac{\log{d}}{d}\alpha\Big)+o(\alpha),$
@@ 2.5.3. The extrapolation
exhibits a monomial in $V(\mathbf{z})^{M}F(\varphi(\mathbf{z}))$ of the minimal order $\beta+M\binom{d}{2}$; for $\big(n_{1}+(d-1)M,n_{2}+(d-2)M,\ldots,n_{d}\big)$ has the strictly highest lexicographical ordering across all monomials of degree $\beta+M\binom{d}{2}$ in $V(\mathbf{z})^{M}F(\varphi(\mathbf{z}))$.
@@
(2.5.23) $\exists\,\mathbf{n}\in\mathbf{N}_{0}^{d}\quad:\quad\log{|c(\mathbf{n})|}\geq\beta\log{|\varphi^{\prime}(0)|}\geq\alpha\log{|\varphi^{\prime}(0)|}.$
We get the requisite holonomy rank bound (2.0.3) on combining the degree bound (1) of Lemma 2.1.2 with the Cauchy upper bound (2.5.22) and the Liouville lower bound (2.5.23), and letting firstly $\alpha\to\infty$, then $d\to\infty$, then $\kappa\to 0$, and finally $\varepsilon\to 0$.

### v4

@@ 2.4.1. Lemma on the lexicographically lowest coefficient
Consider a function $G(\mathbf{z})\in\mathbf{C}\llbracket\mathbf{z}\rrbracket\smallsetminus\{0\}$ holomorphic on the closed unit polydisc $\{\mathbf{z}\,:\,\max_{i=1}^{d}|z_{i}|\leq 1\}$, and let $c\,\mathbf{z^{n}}$ be the lexicographically minimal monomial. Then
@@
By assumption, the lexicographically minimal monomial in $H(0,\mathbf{z}^{\prime})\in\mathbf{C}\llbracket\mathbf{z}^{\prime}\rrbracket$ is equal to $c\,\mathbf{z}^{\prime\mathbf{n}^{\prime}}$, where $\mathbf{n}=(n_{1},\mathbf{n}^{\prime})$. Therefore the induction hypothesis gives
@@
Thus, with $c\,\mathbf{z^{n}}$ the lexicographically lowest monomial in $G(\mathbf{z})$, we get from equation (2.4.3) and Lemma 2.1.2 our Cauchy upper bound:
(2.4.8) $\begin{split}\log{|c|}\leq\int_{\mathbf{T}^{d}}\log{|F(\varphi(z_{1}),\ldots,\varphi(z_{d}))|}\,\mu_{\mathrm{Haar}}\\ \leq dD\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}+\kappa C\,\alpha+o(\alpha),\end{split}$
@@
on the split-variables constituents in (2.1.3). Given our normalizations $x(t)\in t+t^{2}\mathbf{Q}\llbracket t\rrbracket$ and $\varphi(z)\in\varphi^{\prime}(0)z+z^{2}\mathbf{C}\llbracket z\rrbracket$, the lexicographically lowest term of $G(\mathbf{z})\in\mathbf{C}\llbracket\mathbf{z}\rrbracket$ is equal to $\varphi^{\prime}(0)^{\beta}$ times the lexicographically lowest term of $F(x(\mathbf{t}))\in\mathbf{Z}\llbracket\mathbf{t}\rrbracket$, where $\beta:=|\mathbf{n}|=n_{1}+\cdots+n_{d}\geq\alpha$ is the common total degree of these lexicographically lowest terms in $G(\mathbf{z})$ and $F(x(\mathbf{t}))$. By (2.4.9), this entails that the nonzero coefficient
@@ 2.5. A second alternative proof
The remainder of § 2 presents our original argument for Theorem 2.0.1, with the thought that it could still be useful for other settings including potential theory (see 2.5.27). Like § 2.2 and unlike § 2.4, it is based on the leading order jet rather than the overall lexicographically lowest monomial in $F(\mathbf{x})$, and on the pointwise Cauchy integral formula instead of on plurisubharmonicity. Contrastingly to both, it employs a cross-variables equidistribution idea.
@@ 2.5.1. Equidistribution
We start out the same way as with Lemma 2.1.2, but now aim to extrapolate based directly on the pointwise Cauchy bound. The key idea here is that upon substituting $x_{j}=\varphi(z_{j})$ into (2.1.3), the $d\to\infty$ equidistribution on the circle of the uniform independent and identically distributed points $z_{1},\ldots,z_{d}$ will normally get the constituent monomials in (2.1.3) to grow at most at the integrated exponential rate of $dD\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}$. The problem with directly applying the Cauchy bound as in [And89, VIII 1.6] is that it involves a pointwise upper bound on the intervening functions $|p(\varphi(\mathbf{z}))^{\mathbf{k}}|$ on the unit polycircle $\mathbf{z}\in\mathbf{T}^{d}$, and while the Monte Carlo heuristic applies on the majority of $\mathbf{T}^{d}$ under $d\to\infty$, with a probability tending to $1$ roughly speaking at a rate exponential in d-d (this follows by Hoeffding’s concentration inequality with (2.5.8) below), the peaks at the biased part of $\mathbf{T}^{d}$ get overwhelmingly large, and a direct extrapolation with (2.1.3) in this way still only leads to a dimension bound with $\sup_{|z|=1}\log{|p\circ\varphi|}$.
@@
The idea for sifting out the equidistributed tuples $(z_{1},\ldots,z_{d})$ is the following. If the points $z_{1},\ldots,z_{d}$ are poorly distributed in the uniform measure of the circle, the quantity $|V(\mathbf{z})|$ is uniformly exponentially small in $d^{2}$ (Lemma 2.5.9 below). This plays off against the $d^{d/2}=\exp(o(d^{2}))$ bound of Lemma 2.5.3 to sift out the equidistributed points in our pointwise upper bound in the Cauchy integral formula when we extrapolate in § 2.4.6 above. Liouville’s Diophantine lower bound still succeeds like in André [And04, §5], thanks to the chain rule and the integrality of the expansion (2.5.2), but at the Cauchy upper bound we are now aided by the fact that $V(\mathbf{z})^{M}$ is extremely small (an exponential in $Md^{2}$, see (2.5.10)) at the peaks of the pointwise Cauchy bound, where the point $(z_{1},\ldots,z_{d})$ is poorly distributed, while still not too large (subexponential in $Md^{2}$, thanks to Lemma 2.5.3) uniformly throughout the whole polycircle $\mathbf{T}^{d}$.
@@
(2.5.8) $D(z_{1},\ldots,z_{d})\leq 3\Big(\frac{1}{K+1}+\sum_{k=1}^{K}\frac{1}{k}\Big|\frac{z_{1}^{k}+\cdots+z_{d}^{k}}{d}\Big|\Big),\quad\textrm{for all }K\in\mathbf{N},$
@@
Thus we introduce another parameter $\varepsilon>0$, which in the end will be let to approach 0 but only after $d\to\infty$, and we divide the points $\mathbf{z}\in\mathbf{T}^{d}$ into two groups according to whether $D(z_{1},\ldots,z_{d})<\varepsilon$ (the well-distributed points) or $D(z_{1}\ldots,z_{d})\geq\varepsilon$ (the poorly distributed points). For the well-distributed group we use Koksma’s inequality (2.5.7), and for the poorly distributed group we take advantage of the overwhelming damping force of the Vandermonde factor.
@@
Since the qualitative result suffices for our purposes here, we give a soft proof based on compactness. The following argument borrows from Bombieri and Gubler’s exposition [BG06, page 103] of Bilu’s equidistribution theorem. The contrapositive of the requisite statement is the existence of an $\varepsilon\in(0,1]$ with
$\liminf_{d\rightarrow\infty}\left\{\inf_{\mathbf{z}\in\mathbf{T}^{d},D(\mathbf{z})\geq\varepsilon}\frac{1}{d^{2}}\sum_{1\leq i<j\leq d}\log{\frac{1}{|z_{i}-z_{j}|}}\right\}\leq 0.$
(If this quantity is strictly positive for all $\varepsilon\in(0,1]$, then define $c(\varepsilon)>0$ to be that quantity.) Hence, arguing for the contradiction, we suppose that there is an $\varepsilon\in(0,1]$ and an infinite sequence $(z_{1}^{(d)},\ldots,z_{d}^{(d)})\in\mathbf{T}^{d}$ such that
(2.5.11) $\lim_{d\to\infty}\left\{\frac{1}{\binom{d}{2}}\sum_{1\leq i<j\leq d}\log{\frac{1}{|z_{i}^{(d)}-z_{j}^{(d)}|}}\right\}\leq 0,$
@@
(2.5.12) $\textrm{for all }d\in\mathbf{N}_{>0},\quad\quad D(z_{1}^{(d)},\ldots,z_{d}^{(d)})\geq\varepsilon.$
@@
across all probability measures $\nu$ supported by $K$. Since $\mathbf{T}$ is invariant under rotation and $\mu_{\mathbf{T}}$ is unique, we have $\mu_{\mathbf{T}}=\mu_{\mathrm{Haar}}$, and since $I(\mu_{\mathrm{Haar}})=0$, but $\mu\neq\mu_{\mathrm{Haar}}$, we have the strict inequality
@@ 2.5.14. Damping the Cauchy estimate
We combine Lemmas 2.5.3 and 2.5.9 for our choice of the damping term $V(\mathbf{z})^{M}$. In the following, all asymptotics are taken under $\alpha\to\infty$ with respect to all other parameters.
@@
$\begin{split}D(z_{1},\ldots,z_{d})<\varepsilon\quad\Longrightarrow\\ \log{|F(\varphi(z_{1}),\ldots,\varphi(z_{d}))|}\leq dD\,\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}+\kappa C\alpha+O_{p,\varphi}(\varepsilon\,dD)+o(\alpha).\end{split}$
@@
On the poorly distributed but exceptional part $D(z_{1},\ldots,z_{d})\geq\varepsilon$, the sum in (2.5.15) can get as large as $d\sup_{|z|=1}\log{|p\circ\varphi|}$. This trivial bound gives, for all $\mathbf{z}\in\mathbf{T}^{d}$:
(2.5.16) $\log{|F(\varphi(z_{1}),\ldots,\varphi(z_{d}))|}\leq dD\sup_{|z|=1}\log^{+}{|p\circ\varphi|}+\kappa C\alpha+o(\alpha).$
@@
for the remainder of the proof of Corollary 2.0.5 (at the end we will firstly take $d\to\infty$, and only then $\varepsilon\to 0$), and we select the Vandermonde exponent
@@
(2.5.20) $\begin{split}\sup_{\mathbf{z}\in\mathbf{T}^{d}:\,D(z_{1},\ldots,z_{d})\leq\varepsilon}\log{|V(\mathbf{z})^{M}F(\varphi(\mathbf{z}))|}\\ \leq dD\,\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}+\kappa C\alpha+O_{p,\varphi}(\varepsilon\alpha)+O_{\varepsilon,p,\varphi}\Big(\frac{\log{d}}{d}\alpha\Big)+o(\alpha).\end{split}$
@@
(2.5.21) $H(\mathbf{z}):=V(\mathbf{z})^{M}F(\varphi(z_{1}),\ldots,\varphi(z_{d}))=:\sum_{\mathbf{n}\in\mathbf{N}^{d}}c(\mathbf{n})\,\mathbf{z^{n}}\in\mathbf{C}\llbracket\mathbf{z}\rrbracket,$
convergent on the closed unit disc $\|\mathbf{z}\|\leq 1$. For each $\mathbf{n}\in\mathbf{N}^{d}$, the $\mathbf{z^{n}}$ coefficient of $H(\mathbf{z})$ is given by the Cauchy integral formula
@@
[…]
@@
(2.5.24) $\begin{split}\log{|c(\mathbf{n})|}\leq dD\,\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}\\ +O(\kappa\alpha)+O_{p,\varphi}(\varepsilon\,\alpha)+O_{\varepsilon,p,\varphi}\Big(\frac{\log{d}}{d}\alpha\Big)+o(\alpha),\end{split}$
@@ 2.5.25. The extrapolation
exhibits a monomial in $V(\mathbf{z})^{M}F(\varphi(\mathbf{z}))$ of the minimal order $\beta+M\binom{d}{2}$, this is because
$\big(n_{1}+(d-1)M,n_{2}+(d-2)M,\ldots,n_{d}\big)$
has the strictly highest lexicographical ordering across all monomials of degree $\beta+M\binom{d}{2}$ in $V(\mathbf{z})^{M}F(\varphi(\mathbf{z}))$.
@@
(2.5.26) $\sup_{\mathbf{n}\in\mathbf{N}^{d}}\left\{\log{|c(\mathbf{n})|}\right\}\geq\beta\log{|\varphi^{\prime}(0)|}\geq\alpha\log{|\varphi^{\prime}(0)|}.$
We get the requisite holonomy rank bound (2.0.3) on combining the degree bound (part (1) of Lemma 2.1.2) with the Cauchy upper bound (2.5.24) and the Liouville lower bound (2.5.26), and letting firstly $\alpha\to\infty$, then $d\to\infty$, then $\kappa\to 0$, and finally $\varepsilon\to 0$.